OpenStudy (anonymous):

solve the system of equations and give the value of x. ok, now for fractions..... x/7+y/21+1... x/2-y/6+0.... A. x=21/2... B.x=7/2 C. infintiely many solutions. D no solutions.

OpenStudy (kropot72):

Your equations should be written as follows: x/7 + y/21 = 1 x/2 - y/6 = 0 Yes???

OpenStudy (kropot72):

@cralls18 Are you there?

OpenStudy (anonymous):

yess sorry

OpenStudy (anonymous):

yes you are correct... thats how its written

OpenStudy (kropot72):

\[\frac{x}{7}+\frac{y}{21}=1...............(1)\] \[\frac{x}{2}-\frac{y}{6}=0.............(2)\] If you multiply both sides of (1) by 7 and multiply both sides of (2) by 2 you will find both of the modified equations begin with x (the x terms are no longer fractions). Can you do this step and post the results. Then I will explain the next step.

OpenStudy (anonymous):

the denominators?

OpenStudy (kropot72):

Multiply both sides of (1) by 7 as follows: \[(7\times \frac{x}{7})+(7\times \frac{y}{21})=1\times 7\] What do you get when you multiply it out?

OpenStudy (kropot72):

\[7\times \frac{x}{7}=\frac{7x}{7}=?\] \[7\times \frac{y}{21}=\frac{7y}{21}=?\] \[1\times 7=?\]

OpenStudy (anonymous):

i apologize, i stepped away for a minute... ok, fractions is my worse subject in math.... so would i multiply 7*7+x? or make 7 a common denominator with 21?

OpenStudy (kropot72):

7x/7 =?

OpenStudy (anonymous):

1

OpenStudy (kropot72):

7/7 = 1 7x/7 = x Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Good :) Then how about 7y/21 = ? (Hint divide the numerator and then the denominator by 7).

OpenStudy (anonymous):

3

OpenStudy (anonymous):

1/3

OpenStudy (kropot72):

7/21 = 1/3 7y/21 = y/3 Do you follow?

OpenStudy (anonymous):

yes sir/ma'am?

OpenStudy (kropot72):

Great! So now our equation (1) when both sides have been multiplied by 7 has become: \[x+\frac{y}{3}=7..............(3)\] And when we multiply both sides of equation (2) by 2 it becomes: \[x-\frac{y}{3}=0.............(4)\] The next step is to subtract equation (4) from equation (3). If we look at the x terms, we have x minus x = 0. Therefore we have eliminated x. If we look at the y terms, we have y/3 minus (-y/3) = can you work this out?

OpenStudy (anonymous):

y/0.. y--y would be +... 3-3=0.. right?

OpenStudy (anonymous):

so the answer to the question is no solution?

OpenStudy (anonymous):

or infinitely many solutions?

OpenStudy (anonymous):

no it has solution

OpenStudy (kropot72):

This problem does have a solution. \[\frac{y}{3}-(-\frac{y}{3})=\frac{y+y}{3}=\frac{2y}{3}\] Remembering that (a/b) - (-a/b) = (a + a)/b + * + = + - * - = + + * - = - - * + = -

OpenStudy (kropot72):

So the result of subtracting (4) from (3) is as follows: \[\frac{2y}{3}=7\] \[y=\frac{21}{2}\] If this value for y is substituted into (3) we get: \[x+\frac{21}{2\times 3}=7\] \[x=7-\frac{7}{2}=you\ can\ calculate\]

OpenStudy (kropot72):

Choose 2 as a common denominator: \[x=\frac{(7\times 2)-7}{2}=\frac{14-7}{2}=?\]